# Three-Stream Temporal-Shift Attention Network Based on Self-Knowledge Distillation for Micro-Expression Recognition

### v2

@@ Abstract
Micro-expressions are subtle facial movements that occur spontaneously when people try to conceal real emotions. Micro-expression recognition is crucial in many fields, including criminal analysis and psychotherapy. However, micro-expression recognition is challenging since micro-expressions have low intensity and public datasets are small in size. To this end, a three-stream temporal-shift attention network based on self-knowledge distillation called SKD-TSTSAN is proposed in this paper. Firstly, to address the low intensity of muscle movements, we utilize learning-based motion magnification modules to enhance the intensity of muscle movements. Secondly, we employ efficient channel attention modules in the local-spatial stream to make the network focus on facial regions that are highly relevant to micro-expressions. In addition, temporal shift modules are used in the dynamic-temporal stream, which enables temporal modeling with no additional parameters by mixing motion information from two different temporal domains. Furthermore, we introduce self-knowledge distillation into the micro-expression recognition task by introducing auxiliary classifiers and using the deepest section of the network for supervision, encouraging all blocks to fully explore the features of the training set. Finally, extensive experiments are conducted on four public datasets: CASME II, SAMM, MMEW, and CAS(ME)3. The experimental results demonstrate that our SKD-TSTSAN outperforms other existing methods and achieves new state-of-the-art performance. Our code will be available at https://github.com/GuanghaoZhu663/SKD-TSTSAN.
@@ 1 Introduction
Early MER methods are based on manual feature extraction, such as Local Binary Pattern histograms from Three Orthogonal Planes (LBP-TOP) (Pfister et al.,, 2011) and Main Directional Mean Optical-flow (MDMO) (Liu et al.,, 2015). These methods largely rely on manually designed extractors and complicated parameter adjustments. The extracted features are mainly low-level features (Li et al., 2022c, ), therefore it is difficult to recognize the subtle movements in MEs with high accuracy. With the proposal of a series of ME datasets (Yan et al.,, 2014; Davison et al.,, 2016; Ben et al.,, 2021; Li et al., 2022b, ) and the development of deep learning, researchers began to use CNN to extract spatial and temporal features of MEs. For example, Song et al., (2019) proposed a three-stream convolutional neural network (TSCNN) to learn the spatiotemporal information of MEs. Xia and Wang, (2021) trained two dual-stream baseline models called MiNet and MaNet using the ME dataset and macro-expression dataset. By introducing two auxiliary tasks between the two models, MiNet can effectively extract both the dynamic patterns of the muscle motion and the static texture of the face. Although these methods can extract ME features and achieve promising results, there are still two major challenges in MER: the low intensity of MEs and overfitting on small ME datasets (Ben et al.,, 2021). Since MEs have low intensity, networks need to have strong representation ability, which generally means more convolutional layers and computational costs. However, complex models usually overfit on small-scale ME datasets. Therefore, in order to mine subtle ME motions from small ME datasets, more effective blocks need be introduced.
@@
To address the above challenges, we propose a three-stream temporal-shift attention network based on self-knowledge distillation called SKD-TSTSAN in this paper. Firstly, we use the TSCNN (Song et al.,, 2019) as our baseline, which consists of three streams, namely static-spatial stream (S-stream), local-spatial stream (L-stream), and dynamic-temporal stream (T-stream). The inputs of the three streams are the whole facial image, the concatenation of $n\times n$ localized facial images, and optical flow images, respectively. The three recognition streams are finally fused into a fully connected layer for MER tasks. Targeting the low intensity of MEs, we add pre-trained magnification networks (Oh et al.,, 2018) to the head of each convolutional stream based on transfer learning. For an end-to-end framework, a magnification network without an encoder is employed. Since ME only appears in local regions of the face, we incorporate efficient channel attention (ECA) modules (Wang et al.,, 2020) into the L-stream to encourage the network to focus on significant local regions.
@@
Compared with 2D CNNs, 3D CNNs can mine the spatiotemporal information of videos more effectively (Li et al., 2022c, ). However, 3D CNNs generally have a large number of parameters and are easily overfitted on small-scale ME datasets (Li et al., 2022c, ). To capture temporal relations in ME sequences and maintain the complexity of 2D CNNs, we introduce temporal shift modules (TSMs) (Lin et al.,, 2019) in T-stream. In TSM, part of the optical flow features’ channels are shifted along the temporal dimension to realize the information exchange. Moreover, to further enhance the performance of the lightweight network, the macro-expression dataset CK+ (Lucey et al.,, 2010) is employed for pre-training, and self-knowledge distillation (SKD) (Zhang et al.,, 2019) is incorporated. Additional convolutional and fully connected layers are added to form multiple auxiliary classifiers (ACs). By using labels and the model’s final output to train the ACs, we enable all of the model’s blocks to fully explore the training set’s features.
@@
    A motion magnification network is employed to increase the intensity of ME muscle motion, making the features of apex frames and optical flow fields more obvious.
@@
    ECA modules are introduced in the local-spatial stream to ignore redundant information and focus on more meaningful channels. TSMs are added in the dynamic-temporal stream, which achieves information exchange in the temporal dimension with no additional parameters.
@@
    We explore the effectiveness of self-knowledge distillation on the MER tasks, introducing auxiliary classifiers in the network to enable comprehensive learning of features from the training set.
@@
    Extensive experimental results on four ME datasets indicate that SKD-TSTSAN achieves state-of-the-art (SOTA) performance. Furthermore, ablation studies show the effectiveness of each component.
@@ 3.1 Motion Magnification Module
Due to the low intensity of ME, we enhance muscle movements based on transfer learning using the motion magnification network proposed by Oh et al., (2018). To achieve an end-to-end framework and avoid noise in magnified images reconstructed in the decoder, our motion magnification module does not use a decoder. Given that each stream in our three-stream network has different inputs, we add separate motion magnification modules for each stream. Specifically, the input to the motion magnification network includes two frames with small motion displacement. For an ME sequence, the onset frame is the first frame where the ME begins, and the apex frame is the frame where the facial expression reaches maximum intensity, conveying significant emotional information. Therefore, the onset and apex frames are employed as inputs for the motion magnification module of the S-stream, and the concatenation of their $n\times n$ grid divisions as inputs for the L-stream’s motion magnification module. Particularly, since the T-stream extracts temporal information from optical flow fields, we utilize the optical flow corresponding to no motion and the optical flow between the onset and apex frames (or: between the apex and offset frames) as inputs for the motion magnification module in the T-stream.
@@ 3.3 Temporal Shift Module
As shown in Fig. 4(a), in the original TSCNN, the input to the T-stream is the concatenation of horizontal and vertical optical flows between the onset frame, apex frame, and offset frame pairs. It can be seen that 2D CNN without TSM have difficulty in capturing temporal relationships in forward propagation. To introduce TSM, we separate the inputs of the T-stream along the temporal dimension, i.e., the optical flow maps between the onset frame and the apex frame and between the apex frame and the offset frame, containing motion information from two different temporal domains. As illustrated in Fig. 4(b), TSMs are added before all convolutional layers in the intermediate layers of the T-stream to mix information from the two different temporal domains, enhancing the model’s temporal modeling capability. When the ratio of shifted channels in TSM is too big, it may affect the learning of spatial features; if too small, the temporal modeling capability is insufficient to handle complex temporal relationships (Lin et al.,, 2019). Therefore, we shift 1/8 of the channels in the TSM.
@@ 3.4 Self-Knowledge Distillation
where $q_{i}^{c}$ indicates the class probability of the class ${{i}_{th}}$ for the classifier ${{\theta}_{c}}$. $T$ denotes the distillation temperature, where a large $T$ is used to soften the class probability distribution.
@@ 4.1 Datasets
We evaluate our SKD-TSTSAN on four spontaneous ME datasets: the CASME II (Yan et al.,, 2014), SAMM (Davison et al.,, 2016), MMEW (Ben et al.,, 2021), and CAS(ME)3 (Li et al., 2022b, ) databases. The details of these datasets are introduced below.
@@
    The CASME II dataset includes 255 ME samples from 26 participants, which are recorded at 200 fps with a facial resolution of $280\times 340$. These samples are divided into seven categories: Happiness (32), Disgust (63), Repression (27), Surprise (25), Sadness (7), Fear (2), and Others (99), where the numbers in brackets represent the number of corresponding MEs. Due to the small number of samples in the “Sadness” and “Fear” categories, each having fewer than 10 samples, the CASME II dataset is usually used to study a 5-category MER task involving the remaining 5 categories. Additionally, some studies classify emotion categories into 3 categories: the “Positive” emotion category contains “Happiness” class, the “Negative” category contains “repression” and “disgust” classes, and the “Surprise” category includes only the “Surprise” class. Our experiments include both 3- and 5-category tasks on the CASME II dataset.
@@
    The SAMM dataset includes 159 ME samples from 29 participants, recorded at a frame rate of 200 fps and a facial resolution of $400\times 400$. Eight emotion categories are labeled in SAMM: Anger (57), Happiness (26), Surprise (15), Contempt (12), Disgust (9), Fear (8), Sadness (6), and Others (26). In our experiments, we considered the 3-category task. The 3-category task includes a “Positive” category that includes “Happiness” class, a “Negative” category that includes “Anger”, “Disgust”, “Contempt”, “Sadness”, “Fear” classes, and a “Surprise” category that included only “Surprise” class.
@@
    The MMEW dataset includes 300 ME samples and 900 macro-expression samples from 36 participants. They are recorded using a frame rate of 90 fps with a facial resolution of $400\times 400$. The ME samples are divided into seven categories: Happiness (36), Disgust (72), Surprise (89), Fear (16), Sadness (13), Anger (8) and Others (66). In our experiments, we use four and six emotion categories respectively. For the 6-class MER task on the MMEW dataset, the “Anger” category is excluded. When performing the 4-class MER task, we select four categories that are close in number: Happiness, Disgust, Surprise, and Others.
@@
    The CAS(ME)3 dataset is a large-scale spontaneous ME dataset, which provides about 80 hours of video at a frame rate of 30 fps and offers depth information as an additional modality. According to the latest version of the dataset annotations, there are 860 manually labeled MEs, divided into seven categories: Anger (64), Disgust (250), Fear (86), Happy (55), Others (161), Sad (57), and Surprise (187). In our experiments, classification tasks are conducted with 3,4, and 7 emotion categories. For the 3-category classification, we group negative emotions (“Anger”, “Disgust”, “Fear”, and “Sad”) into a “Negative” category, “Happy” into a “Positive” category, and “Surprise” into the third category. In the 4-category classification, we add the “Others” category. The 7-category classification uses all the official emotion categories.
@@ 4.2 Data Pre-Processing and Implementation Details
As the proposed SKD-TSTSAN requires knowledge transfer from macro-expression data, we first pre-train it on the CK+ dataset (Lucey et al.,, 2010), which includes seven expressions: Happiness, Anger, Disgust, Contempt, Fear, Sadness, and Surprise. The CK+ dataset, being a macro-expression dataset, lacks annotations like the apex frames present in ME datasets. Therefore, to pre-train the network using this dataset, we designate the last frame with the most obvious muscle movement of each sample as the apex frame, and the first frame as both the onset and offset frames.
@@
We implement our SKD-TSTSAN on an NVIDIA-RTX-4090 GPU. The Adam optimizer is utilized with learning rate, betas and weight decay set to 1e-3, (0.9, 0.99) , 5e-4, respectively. Loss weights ${{\lambda}_{1}}=0.1$, ${{\lambda}_{2}}=1e-6$ and distillation temperature $T=3$ are chosen as default values. In the multi-class focal loss, the focusing parameter $\gamma$ is set to 2, and the weighting factor ${{\alpha}_{k}}$ is positively correlated with the inverse of the sample size for each category. In our motion magnification module, we set the amplification factor to 2. During the pre-training stage on the CK+ dataset, we configure the training iterations to 10k and the batch size to 16. When training on the CASME II, SAMM, and MMEW datasets, the batch size is set to 16, while for the CAS(ME)3 dataset, it is set to 32. Training iterations on the ME dataset are set to 20k and early stopping is used to alleviate overfitting.

### v3

@@ Abstract
Micro-expressions are subtle facial movements that occur spontaneously when people try to conceal real emotions. Micro-expression recognition is crucial in many fields, including criminal analysis and psychotherapy. However, micro-expression recognition is challenging since micro-expressions have low intensity and public datasets are small in size. To this end, a three-stream temporal-shift attention network based on self-knowledge distillation is proposed in this paper. Firstly, to address the low intensity of muscle movements, we utilize learning-based motion magnification modules to enhance the intensity of muscle movements. Secondly, we employ efficient channel attention modules in the local-spatial stream to make the network focus on facial regions that are highly relevant to micro-expressions. In addition, temporal shift modules are used in the dynamic-temporal stream, which enables temporal modeling with no additional parameters by mixing motion information from two different temporal domains. Furthermore, we introduce self-knowledge distillation into the micro-expression recognition task by introducing auxiliary classifiers and using the deepest section of the network for supervision, encouraging all blocks to fully explore the features of the training set. Finally, extensive experiments are conducted on five publicly available micro-expression datasets. The experimental results demonstrate that our network outperforms other existing methods and achieves new state-of-the-art performance. Our code is available at https://github.com/GuanghaoZhu663/SKD-TSTSAN.
@@ 1 Introduction
Early MER methods are based on manual feature extraction, such as Local Binary Pattern histograms from Three Orthogonal Planes (LBP-TOP) (Pfister et al.,, 2011) and Main Directional Mean Optical-flow (MDMO) (Liu et al.,, 2015). These methods largely rely on manually designed extractors and complicated parameter adjustments. The extracted features are mainly low-level features (Li et al., 2022c, ), therefore it is difficult to recognize the subtle movements in MEs with high accuracy. With the proposal of a series of ME datasets (Yan et al.,, 2014; Davison et al.,, 2016; Li et al.,, 2013; Ben et al.,, 2021; Li et al., 2022b, ) and the development of deep learning, researchers began to use CNN to extract spatial and temporal features of MEs. For example, Song et al., (2019) proposed a three-stream convolutional neural network (TSCNN) to learn the spatiotemporal information of MEs. Xia and Wang, (2021) trained two dual-stream baseline models called MiNet and MaNet using the ME dataset and macro-expression dataset. By introducing two auxiliary tasks between the two models, MiNet can effectively extract both the dynamic patterns of the muscle motion and the static texture of the face. Although these methods can extract ME features and achieve promising results, there are still two major challenges in MER: the low intensity of MEs and overfitting on small ME datasets (Ben et al.,, 2021). Since MEs have low intensity, networks need to have strong representation ability, which generally means more convolutional layers and computational costs. However, complex models usually overfit on small-scale ME datasets. Therefore, in order to mine subtle ME motions from small ME datasets, more effective blocks need be introduced.
@@
To address the above challenges, we propose a three-stream temporal-shift attention network based on self-knowledge distillation called SKD-TSTSAN in this paper. Firstly, we use the TSCNN (Song et al.,, 2019) as our baseline, which consists of three streams, namely static-spatial stream (S-stream), local-spatial stream (L-stream), and dynamic-temporal stream (T-stream). The inputs of the three streams are the whole facial image, the concatenation of $n\times n$ localized facial images, and optical flow images, respectively. The three recognition streams are finally fused into a fully connected layer for MER tasks.
To tackle the low intensity of MEs, we introduce a motion magnification network (Oh et al.,, 2018) at the head of each convolutional stream. This module enhances subtle muscle motions, making the features of apex frames and optical flow fields more discriminative. Since MEs typically appear in localized facial regions, such as the mouth and eyebrows, the input to the L-stream often contains redundant information. To address this, we incorporate efficient channel attention (ECA) modules (Wang et al.,, 2020) into the L-stream, enabling the network to focus on meaningful local regions while ignoring redundant information.
@@
Compared with 2D CNNs, 3D CNNs can mine the spatiotemporal information of videos more effectively (Li et al., 2022c, ). However, 3D CNNs generally have a large number of parameters and are easily overfitted on small-scale ME datasets (Li et al., 2022c, ). To effectively model temporal dynamics without the computational overhead of 3D CNNs, we introduce temporal shift modules (TSMs) (Lin et al.,, 2019) in T-stream. These modules enable efficient information exchange along the temporal dimension without introducing additional parameters. To further enhance the performance of our lightweight network, the macro-expression dataset CK+ (Lucey et al.,, 2010) is employed for pre-training. Moreover, we employ self-knowledge distillation (SKD) (Zhang et al.,, 2019). Traditional knowledge distillation requires a complex teacher model to guide a compact student model (Hinton et al.,, 2015), but designing an effective over-parameterized teacher model is challenging on small-scale micro-expression datasets. In SKD, we introduce additional convolutional and fully connected layers to form multiple auxiliary classifiers (ACs). By training ACs using both labels and the model’s final output, we ensure that all layers of the network comprehensively learn the training set’s features.
@@
    A motion magnification network is employed to enhance the intensity of ME muscle motion, making the features of apex frames and optical flow fields more discriminative.
@@
    ECA modules are introduced in the local-spatial stream to focus on meaningful facial regions while ignoring redundant information.
@@
    TSMs are integrated into the dynamic-temporal stream, enabling efficient temporal modeling without additional parameters.
@@
    We explore the effectiveness of self-knowledge distillation on the MER tasks, introducing auxiliary classifiers in the network to enable comprehensive feature learning from the training set.
  • 5.
    Extensive experiments on five ME datasets demonstrate that SKD-TSTSAN achieves state-of-the-art (SOTA) performance. Ablation studies further validate the effectiveness of each module.
@@ 3.1 Motion Magnification Module
Due to the low intensity of ME, we enhance muscle movements using the motion magnification network proposed by Oh et al., (2018). To achieve an end-to-end framework and avoid noise in magnified images reconstructed in the decoder, our motion magnification module does not use a decoder. Given that each stream in our three-stream network has different inputs, we add separate motion magnification modules for each stream. Specifically, the input to the motion magnification network includes two frames with small motion displacement. For an ME sequence, the onset frame is the first frame where the ME begins, and the apex frame is the frame where the facial expression reaches maximum intensity, conveying significant emotional information. Therefore, the onset and apex frames are employed as inputs for the motion magnification module of the S-stream, and the concatenation of their $n\times n$ grid divisions as inputs for the L-stream’s motion magnification module. Particularly, since the T-stream extracts temporal information from optical flow fields, we utilize the optical flow corresponding to no motion and the optical flow between the onset and apex frames (or: between the apex and offset frames) as inputs for the motion magnification module in the T-stream.
@@ 3.3 Temporal Shift Module
As shown in Fig. 4(a), in the original TSCNN, the input to the T-stream is the concatenation of horizontal and vertical optical flows between the onset frame, apex frame, and offset frame pairs. It can be seen that 2D CNN without TSM have difficulty in capturing temporal relationships in forward propagation. To introduce TSM, we separate the inputs of the T-stream along the temporal dimension, i.e., the optical flow maps between the onset frame and the apex frame and between the apex frame and the offset frame, containing motion information from two different temporal domains. As illustrated in Fig. 4(b), TSMs are added before all convolutional layers in the intermediate layers of the T-stream to mix information from the two different temporal domains, enhancing the model’s temporal modeling capability. When the ratio of shifted channels in TSM is too big, it may affect the learning of spatial features; if too small, the temporal modeling capability is insufficient to handle complex temporal relationships (Lin et al.,, 2019). Therefore, in the TSM, we shift the first 1/8 of the channels one frame forward, shift the following 1/8 of the channels one frame backward, and leave the remaining 3/4 unchanged.
@@ 3.4 Self-Knowledge Distillation
where $q_{i}^{c}$ indicates the class probability of the ${{i}_{th}}$ class for the classifier ${{\theta}_{c}}$. $T$ denotes the distillation temperature, where a large $T$ is used to soften the class probability distribution.
@@ 4.1 Datasets
We evaluate our SKD-TSTSAN on five spontaneous ME datasets: the CASME II (Yan et al.,, 2014), SAMM (Davison et al.,, 2016), SMIC (Li et al.,, 2013), MMEW (Ben et al.,, 2021), and CAS(ME)3 (Li et al., 2022b, ) databases. The details of these datasets are introduced below.
@@
    The CASME II dataset includes 255 ME samples from 26 participants, which are recorded at 200 fps with a facial resolution of $280\times 340$. These samples are divided into seven categories: Happiness (32), Disgust (61), Repression (27), Surprise (25), Sadness (7), Fear (2), and Others (99), where the numbers in brackets represent the number of corresponding MEs. Due to the small number of samples in the “Sadness” and “Fear” categories, each having fewer than 10 samples, the CASME II dataset is usually used to study a 5-category MER task involving the remaining 5 categories. Additionally, some studies classify emotion categories into 3 categories: the “Positive” emotion category contains “Happiness” class, the “Negative” category contains “repression” and “disgust” classes, and the “Surprise” category includes only the “Surprise” class.
@@
    The SAMM dataset includes 159 ME samples from 29 participants, recorded at a frame rate of 200 fps and a facial resolution of $400\times 400$. Eight emotion categories are labeled in SAMM: Anger (57), Happiness (26), Surprise (15), Contempt (12), Disgust (9), Fear (8), Sadness (6), and Others (26). This dataset is typically divided into three categories: a ”Positive” category, which includes the ”Happiness” class; a ”Negative” category, comprising the ”Anger”, ”Disgust”, ”Contempt”, ”Sadness”, and ”Fear” classes; and a ”Surprise” category, which consists solely of the ”Surprise” class.
@@
    The SMIC dataset provides three subsets recorded with different types of cameras: HS (high-speed camera), VIS (visible light camera), and NIR (near-infrared camera). In our experiments, we adopted the HS subset, which was collected using a high-speed camera at a frame rate of 100 fps. This subset consists of 164 ME samples from 16 participants, with a facial resolution of $190\times 230$. The emotion categories in SMIC are labeled as Negative (70), Positive (51), and Surprise (43). Since the SMIC dataset does not provide labels for apex frames, we utilized the apex frame labels provided by the CapsuleNet method (Van Quang et al.,, 2019).
@@
    The MMEW dataset includes 300 ME samples and 900 macro-expression samples from 36 participants. They are recorded using a frame rate of 90 fps with a facial resolution of $400\times 400$. The ME samples are divided into seven categories: Happiness (36), Disgust (72), Surprise (89), Fear (16), Sadness (13), Anger (8) and Others (66). The MMEW dataset is typically divided into four and six emotion categories. For the 6-class MER task, the ”Anger” category is excluded. In the 4-class MER task, four categories with similar sample sizes are selected: Happiness, Disgust, Surprise, and Others.
  5. 5.
    The CAS(ME)3 dataset is a large-scale spontaneous ME dataset, which provides about 80 hours of video at a frame rate of 30 fps and offers depth information as an additional modality. According to the latest version of the dataset annotations, there are 860 manually labeled MEs, divided into seven categories: Anger (64), Disgust (250), Fear (86), Happy (55), Others (161), Sad (57), and Surprise (187). This dataset is typically divided into 3, 4, and 7 emotion categories. For the 3-category classification, negative emotions such as ”Anger”, ”Disgust”, ”Fear”, and ”Sad” are grouped into a ”Negative” category, ”Happy” is assigned to a ”Positive” category, and ”Surprise” forms the third category. In the 4-category classification, an additional ”Others” category is included. The 7-category classification utilizes all the official emotion categories.
To comprehensively evaluate the effectiveness of our SKD-TSTSAN in the field of MER, we conducted both composite dataset evaluation (See et al.,, 2019) and single dataset evaluation in our experiments. For the composite dataset evaluation, CASME II, SAMM, and SMIC were merged into a 3-category composite dataset, where the MER task utilized the shared emotion labels across these datasets: Positive (109 samples), Negative (250 samples), and Surprise (83 samples). In the single dataset evaluation, experiments were performed on individual datasets. Specifically, we conducted 5-class MER on the CASME II dataset, 4-class and 6-class MER on the MMEW dataset, and 3-class, 4-class, and 7-class MER on the CAS(ME)3 dataset.
@@ 4.2 Data Pre-Processing and Implementation Details
As the proposed SKD-TSTSAN requires knowledge transfer from macro-expression data, we first pre-train it on the CK+ dataset (Lucey et al.,, 2010), which includes seven expressions: Happiness, Anger, Disgust, Contempt, Fear, Sadness, and Surprise. The CK+ dataset, being a macro-expression dataset, lacks annotations like the apex frames present in ME datasets. Therefore, to pre-train the network using this dataset, we designate the last frame with the most obvious muscle movement of each sample as the apex frame, and the first frame as both the onset and offset frames. After pre-training, we employ transfer learning to transfer the knowledge learned from the macro-expression dataset to the MER task. Specifically, we remove the final linear layer and use the remaining weights as the initial weights for training on the micro-expression dataset. By transferring knowledge from macro-expressions to micro-expressions, the model benefits from the rich information in macro-expression data while adapting to the subtle and rapid dynamics of micro-expressions.
@@
We implement our SKD-TSTSAN on an NVIDIA-RTX-4090 GPU. The Adam optimizer is utilized with learning rate, betas and weight decay set to 1e-3, (0.9, 0.99) , 5e-4, respectively. Loss weights ${{\lambda}_{1}}=0.1$, ${{\lambda}_{2}}=1e-6$ and distillation temperature $T=3$ are chosen as default values. In the multi-class focal loss, the focusing parameter $\gamma$ is set to 2, and the weighting factor ${{\alpha}_{k}}$ is positively correlated with the inverse of the sample size for each category. In our motion magnification module, we set the amplification factor to 2. During the pre-training stage on the CK+ dataset, we configure the training iterations to 10k and the batch size to 16. When training on the CASME II, MMEW, and composite datasets, the batch size is set to 16, while for the CAS(ME)3 dataset, it is set to 32. Training iterations on the ME dataset are set to 20k and early stopping is used to alleviate overfitting.